### v3

@@ Abstract.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations. We also find that mini-batch training techniques converge to similar or often higher accuracy values as full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy. Our work highlights the importance of comparing systems across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for each training method separately.
@@ 1. Introduction
The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019). Mini-batch GNN training introduces additional variability in the gradient estimation, even compared to regular mini-batch DNN training. Specifically, in regular DNN training, the stochasticity in gradient estimation primarily arises from using mini-batches, which are random subsets of the training data, in each iteration. In mini-batch GNN training, the graph sampling used to form micro-batches is an additional source of stochasticity beyond the sampling of vertices in the mini-batch, because the training algorithm is not exposed to all the required dependent data in the graph.
@@
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes. For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., [n.d.]b). For mini-batch training, we consider DGL (Wang et al., 2019), DistDGL (Zheng et al., 2020), and Quiver (Tan et al., 2023), and three sampling algorithms: Neighborhood Sampling (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN models, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, showing that the latter has a larger epoch time  (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. We report our performance results in Section 4 and support and generalize our empirical observations with an analysis in Section 6.
@@
Overall, our work highlights that existing algorithmic optimizations such as sampling or asynchrony can improve efficiency with no or minimal accuracy degradation, representing a promising research direction to achieve further performance and accuracy improvements. From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes and it also indicates a set of principles on how to perform this comparison.
@@ 2.1. Graph Neural Networks
Proposes System/paper Main FG vs. MB Claims Experimental evidence Year Open Source Evaluation Setup
@@
G3 (Wan et al., 2023b) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023c) MB information loss 2023 Multi-host multi-GPU
@@
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. In this paper, we focus our evaluation on GPU-based training on multi-GPU systems, which is a common configuration for scalable GNN training. For a comprehensive and detailed discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems. The key idea of full-graph training is to execute the GNN layers for all vertices in the graph, as shown in Eqn. 1. To scale to large graphs, distributed full-graph training systems partition the graph into multiple subgraphs such that each can fit into the memory of one GPU. They use a model-parallel approach to train across GPUs that requires exchanging vertex features across partitions. The main focus of work on full-graph training systems has been on reducing the cost of vertex feature communication. We now discuss existing full-graph GNN training systems work in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
Some full-graph training systems proposed optimizations that focused on the management of training data and do not introduce approximations that can potentially impact accuracy. If there are not enough GPUs to store each partition in a different GPU, it is necessary to transfer data between the host memory and GPU memory. Prior work such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023b), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication. When there are enough GPUs to distribute all the partitions of the graph in GPU memory, training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023b; Cai et al., 2021; Tripathy et al., 2020).
Some recent works improve performance by further introducing algorithmic optimizations that can impact accuracy, unlike the work we described previously. Sancus (Peng et al., 2022), PipeGCN (Wan et al., 2022b), and GNNPipe (Chen et al., 2023) introduce asynchrony and allow GPUs to operate on stale vertex features. This enables GPUs to overlap communication with computation instead of having to wait at each layer for fresh vertex features coming from other GPUs. An alternative research direction has been to use sampling. BNS-GCN samples boundary nodes with edges across partitions at each epoch and exchanges vertex features only for those vertices (Wan et al., 2022a). ADGNN proposes an aggregation-difference sampling algorithm (Song et al., 2023a). Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023c).
@@ 2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of work on GNN training do not compare or evaluate their systems against systems using a different training approach and training pipeline (full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). Some prior work focuses only on the data-parallel mini-batching pipeline of Figure 1(b) and compares the performance of different low-level data management optimizations on top of the same underlying system (Yuan et al., 2024). In the following, we review 33 papers on GNN training that make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
@@
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory, and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of the scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time-to-accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023b, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
@@ 3. Experimental methodology
For full-graph training, in the training data management dimension of Figure 2, our evaluation only considers the more favorable situation where the entire training data is fully cached in the memory of the GPUs since they can entirely avoid the overhead of GPU-host communication. All systems adopt a pipeline similar to the one of Figure 1(a). The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023a). In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023a), BNS-GCN for sampling (Wan et al., 2022a, [n.d.]a), and AdaQP for message quantization (Wan et al., 2023c, [n.d.]b).
For mini-batch training, we consider three baselines: DGL (Wang et al., 2019; dgl, 2023), its distributed version DistDGL (Zheng et al., 2020; dgl, 2023), and Quiver (Tan et al., 2023; Quiver, 2023). These cover the system-level optimizations discussed in Section 2.2 (see Figure 2). All systems adopt a pipeline similar to the one of Figure 1(b). In terms of sampling implementations, DistDGL uses CPU-based sampling, while both DGL and Quiver support GPU-based sampling. DistDGL partitions the training dataset across multiple hosts. In terms of GPU caching, DGL only supports local GPU caching while Quiver also supports distributed GPU caching. Quiver uses distributed input feature caching for the Orkut and Papers100M datasets, which are too large for local caching. Mini-batch training can use different sampling algorithms to select a subset of the k-hop neighbors of the vertices in the mini-batch. We consider the standard Neighborhood Sampling (NS) algorithm, which is available for both DGL and Quiver, ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
@@ 4.2. Scalability
Multi-host scalability (NVLink). We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has 4 GPUs. For this experiment, we replace DGL with DistDGL, its distributed version. We report the hyperparameters in the appendix in table 15. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We run Quiver with data parallelism across multiple hosts.
@@ 5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target test accuracy, it could still be preferable to use it if it can converge to a higher test accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
@@
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the test best accuracy we can achieve using full-graph and mini-batch training? Q4) What test accuracy can we achieve when using a training method with the best hyperparameters found for the other method? Different training systems have different optimizations, so Q5: What is the impact of the optimizations on accuracy?
@@
Our evaluation of GAT and GCN is reported on the appendix in table 9 and table 10 which leads to similar conclusions. For GAT, the largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for arxiv and $1.9\%$ in favor of FG-search for pubmed. For GCN, it is $1.36\%$ in favor of FG-search for reddit and $5.5\%$ in favor of MB-search for pubmed.
@@
Impact of optimizations on accuracy. Many performance optimizations we consider can result in a different accuracy than the vanilla methods. This section answers the question Q5: What is the impact of the optimizations on accuracy? In this analysis, we use the best-found hyperparameters for each system class: FG-Search and MB-Search models for full-graph and mini-batch training systems, respectively. For full-graph training, we consider the impact of asynchrony (PipeGCN), sampling (BNS-GCN), and message quantization (AdaQP). For mini-batch training, besides Neighborhood Sampling (NS), we also consider the ClusterGCN and GraphSaint sampling algorithms.

### v4

@@ Abstract.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations. We also find that mini-batch training techniques converge to similar to or often higher accuracy values than full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy. Our work highlights the importance of comparing systems across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for each training method separately.
@@ 1. Introduction
The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019). Mini-batch GNN training introduces additional variability in the gradient estimation, even compared to regular mini-batch DNN training. Specifically, in regular DNN training, the stochasticity in gradient estimation primarily arises from using mini-batches, which are random subsets of the training data, in each iteration. In mini-batch GNN training, the graph sampling used to form micro-batches is an additional source of stochasticity beyond the sampling of vertices in the mini-batch, because the training algorithm is not exposed to all the required dependent data in the graph.
@@
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes. For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., 2024). For mini-batch training, we consider DGL (Wang et al., 2019), DistDGL (Zheng et al., 2020), and Quiver (Tan et al., 2023), and three sampling algorithms: Neighborhood Sampling (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN models, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, and show that mini-batch systems have a larger epoch time  (Wang et al., 2022; Wan et al., 2023c; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. We report our performance results in Section 4 and support and generalize our empirical observations with an analysis in Section 6.
@@
Overall, our work highlights that existing algorithmic optimizations such as sampling or asynchrony can improve efficiency with no or minimal accuracy degradation, representing a promising research direction to achieve further performance and accuracy improvements. From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes and it also indicates a set of principles on how to perform this comparison. The source code, data, and other artifacts used for this evaluation are publicly available (Bajaj et al., 2024).
@@ 2.1. Graph Neural Networks
Proposes System / Paper Main FG vs. MB Claims Experimental Evidence Year Evaluation Setup
@@
G3 (Wan et al., 2023c) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023d) MB information loss 2023 Multi-host multi-GPU
@@
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. In this paper, we focus our evaluation on GPU-based training on multi-GPU systems, which is a common configuration for scalable GNN training. We consider system-level optimizations of the vanilla full-graph and mini-batch training pipelines that improve performance and support any GNN model, as defined in Section 2.1. We show the classes of optimizations we consider and the related baselines in Figure 2. For a comprehensive discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems. The key idea of full-graph training is to execute the GNN layers for all vertices in the graph, as shown in Eqn. 1. To scale to large graphs, distributed full-graph training systems partition the graph into multiple subgraphs such that each can fit into the memory of one GPU. They use a model-parallel approach to train across GPUs that requires exchanging vertex features across partitions. The main focus of work on full-graph training systems has been on reducing the cost of vertex feature communication. We now discuss existing work on full-graph GNN training systems in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
Some full-graph training systems focused on optimizing the management of training data without introducing approximations that can potentially impact accuracy. If there are not enough GPUs to store each partition in a different GPU, it is necessary to transfer data between the host memory and GPU memory. Prior works such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023c), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication. When there are enough GPUs to distribute all the partitions of the graph in GPU memory, these optimizations are not required and training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023c; Cai et al., 2021; Tripathy et al., 2020).
Some recent works improve performance by further introducing algorithmic optimizations that can impact accuracy, unlike the work we described previously. Sancus (Peng et al., 2022), PipeGCN (Wan et al., 2022b), and GNNPipe (Chen et al., 2023) introduce asynchrony and allow GPUs to operate on stale vertex features. This enables GPUs to overlap communication with computation instead of having to wait at each layer for fresh vertex features coming from other GPUs. An alternative research direction has been to use sampling. BNS-GCN samples boundary nodes with edges across partitions at each epoch and exchanges vertex features only for those vertices (Wan et al., 2022a). ADGNN proposes an aggregation-difference sampling algorithm (Song et al., 2023a). Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023d).
@@ 2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of work on GNN training do not compare or evaluate their systems against systems using a different training approach and training pipeline (full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). Some prior work focused on the mini-batch training pipeline of Figure 1(b), evaluating the impact of using different data partitioning algorithms, configurations of the neighborhood sampling algorithm, and data loading optimizations within the same system (Yuan et al., 2024). In the following, we review 33 papers on GNN training that make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
@@
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory, and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of the scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023c; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time-to-accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023c, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
@@ 3. Experimental methodology
For full-graph training, in the training data management dimension of Figure 2, our evaluation only considers the more favorable situation where the entire training data is fully cached in the memory of the GPUs since they can entirely avoid the overhead of GPU-host communication. All systems adopt a pipeline similar to the one of Figure 1(a). The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023b). In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023b), BNS-GCN for sampling (Wan et al., 2022a, 2023a), and AdaQP for message quantization (Wan et al., 2023d, 2024).
For mini-batch training, we consider three baselines: DGL (Wang et al., 2019; dgl, 2023), its distributed version DistDGL (Zheng et al., 2020; dgl, 2023), and Quiver (Tan et al., 2023; Quiver, 2023). These cover the system-level optimizations discussed in Section 2.2 (see Figure 2). All systems adopt a pipeline similar to the one of Figure 1(b). In terms of sampling implementations, DistDGL uses CPU-based sampling, while both DGL and Quiver support GPU-based sampling. DistDGL partitions the training dataset across multiple hosts. All systems use GPU caching whenever possible. DGL only supports local GPU caching while Quiver also supports distributed GPU caching, which is necessary to cache the Orkut and Papers100M datasets. Mini-batch training can use different sampling algorithms to select a subset of the k-hop neighbors of the vertices in the mini-batch. We consider the standard Neighborhood Sampling (NS) algorithm, which is available for both DGL and Quiver, ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
@@ 4.2. Scalability
Multi-host scalability (NVLink). We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has 4 GPUs. For this experiment, we replace DGL with DistDGL, its distributed version. We report the hyperparameters in the in the appendix table 15. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We run Quiver with data parallelism across multiple hosts.
@@ 5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target test accuracy, it could still be preferable to use it if it can converge to a higher test accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
@@
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best test accuracy we can achieve using full-graph and mini-batch training? Q4) What test accuracy can we achieve when using a training method with the best hyperparameters found for the other method? Different training systems have different optimizations (see Figure 2), so Q5: What is the impact of the optimizations on accuracy?
@@
Our evaluation of GAT and GCN is reported in the appendix in table 9 and table 10 which leads to similar conclusions. For GAT, the largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for arxiv and $1.9\%$ in favor of FG-search for pubmed. For GCN, it is $1.36\%$ in favor of FG-search for reddit and $5.5\%$ in favor of MB-search for pubmed.
@@
Impact of optimizations on accuracy. Many performance optimizations we consider can result in a different accuracy than the vanilla methods (see Figure 2). This section answers the question Q5: What is the impact of the optimizations on accuracy? In this analysis, we use the best-found hyperparameters for each system class: FG-Search and MB-Search models for full-graph and mini-batch training systems, respectively. For full-graph training, we consider the impact of asynchrony (PipeGCN), sampling (BNS-GCN), and message quantization (AdaQP). For mini-batch training, besides Neighborhood Sampling (NS), we also consider the ClusterGCN and GraphSaint sampling algorithms.